OpenStudy (anonymous):

When a certain number is increased by 25% its value becomes 388. What will the value be if the original number is decreased by 30%?

OpenStudy (tkhunny):

Please NAME the "certain number".

OpenStudy (anonymous):

They didn't say

OpenStudy (anonymous):

But I think it's 291

OpenStudy (tkhunny):

Seriously? I did not ask what name was given in the problem statement. I asked YOU to name it. Go ahead. Just give it a name.

OpenStudy (anonymous):

x

OpenStudy (tkhunny):

Perfect. Can you write an expression for 25% of x?

OpenStudy (anonymous):

0.25x?

OpenStudy (tkhunny):

Excellent. Can you now write an expression for x increased by 25% of x?

OpenStudy (anonymous):

x+0.25x

OpenStudy (tkhunny):

Superb. And this is the same as x(1 + 0.25) = x(1.25) Do you agree with this assessment?

OpenStudy (anonymous):

Yes

OpenStudy (tkhunny):

Okay, now we have x + 0.25x = 388 or x*(1 + 0.25) = 388 or x*(1.25) = 388 If you agree, can you now solve for x?

OpenStudy (anonymous):

x = 310.40

OpenStudy (tkhunny):

You are awesome. Where does that leave us? What is that number increased by 30% of itself?

OpenStudy (anonymous):

I have no idea

OpenStudy (tkhunny):

Can you write an expression that represents 30% of the value we are calling 'x'?

OpenStudy (anonymous):

Would it be 0.30x ?

OpenStudy (tkhunny):

Perfect. Just like the 25% we did earlier. Now, how about the value, x increased by 30% of itself?

OpenStudy (anonymous):

x+0.30x

OpenStudy (tkhunny):

You have it. x + 0.30x = x(1 + 0.30) = x(1.30) We know the value of x. x = 310.40 What is 310.40 * 1.30?

OpenStudy (anonymous):

403.52

OpenStudy (tkhunny):

That's what I get.

OpenStudy (anonymous):

Hmmm, it's saying that's incorrect

OpenStudy (tkhunny):

Whoops! Another reading of the problem says we DECREASE by 30%, not INCREASE.

OpenStudy (tkhunny):

Start back here... Can you write an expression representing a number we are calling 'x' DECREASED by 30% of its value? Keep in mind that 30% of the value is 0.30x.

OpenStudy (anonymous):

x-0.30x

OpenStudy (tkhunny):

That's it. Can you now solve the problem?

OpenStudy (anonymous):

217.28

OpenStudy (anonymous):

Thanks!

OpenStudy (tkhunny):

NOW we have it. Lesson learned? ALWAYS reread the problem statement before you think you are done. Let's just make sure we are answering the right question!

OpenStudy (whpalmer4):

Good practice to always reread the problem before checking your answer. Many times you'll discover that you've actually answered some other question. And always check the answer!

OpenStudy (tkhunny):

While we're beating the dead horse of philosophy, this is not ALWAYS a good rule. If you are under extreme time pressure, such as GRE, LSAT, GMAT or various competitions and important exams, you simply don't have time to do that. Do the best you can and move on! Under normal circumstances, it IS a GREAT rule to follow.